# BConformeR: A Conformer Based on Mutual Sampling for Unified Prediction of Continuous and Discontinuous Antibody Binding Sites

### v1

@@ 5 Conclusion and Limitation
There are two notable limitations to this study. First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms. Second, despite the size of our collection, antigen families and epitope types are imbalanced, which could bias both training and evaluation.
@@ Acknowledgments
Figure 6: Linear regions and discontinuous fragments. (A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to parts of the two antibody heavy chains; Linear regions are marked with red rectangles, and discontinuous fragments with green. (Right) Residues highlighted with numbers indicate epitopes. (B) A linear region of 7c88. Gap size of up to three non-contacting residues is chosen for a linear region on the basis of the structure of $\alpha$-helices.
@@ Appendix B Linear Regions and Discontinuous Fragments
Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues (Fig. 6A). Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included (Fig. 6B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig. 6A).
@@ C.1 Structure prediction
We initially obtained 1730 antibody–antigen (Ab–Ag) FASTA files from AbEpiTope-1.0. However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3. Each of these 1300 complexes was modeled with 5 seeds, 6 models and 3 recycles, generating 30 structures per complex.
@@ C.2 Epitope statistics
The final dataset consists of 1180 Ab–Ag complexes, which contains a total of 1432 antigen chains and 2320 antibody chains. Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues. After classifying the epitopes into linear and discontinuous types, we found 2388 linear regions comprising 15,886 residues, and 2632 discontinuous fragments comprising 3574 residues.
@@ Appendix D Performance Metrics
The performance metrics used in this study include Precision, Recall, F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
$\text{Precision}=\frac{TP}{TP+FP},$ (9)
@@
where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means. The relationships between these metrics are given as:
$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$ (10)
@@ E.2 Baseline methods
Thresholds for all baselines were adjusted to reflect their best scoring behavior, ensuring a fair comparison with BConformeR in the blind test. We observed that SEPPA-3.0 performs better under a low decision threshold (around 0.1). For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351. Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.2 led to better balance between precision and recall on our test set.
@@ F.1 ViT
ViT-9 1536 1024 1 9 12 2.0
ViT-12 1536 1024 1 12 12 2.0
Table 4: ViT-1D variants used in ablation study.
We conducted ablation studies on two variants of the ViT-1D model differing in depth: ViT-9 and ViT-12. As shown in Table 4, both models share the same embedding dimension (1536), input sequence length (1024), patch size (1), number of attention heads (12), and MLP expansion ratio (2.0). The only difference is the number of transformer encoder layers.
F.2 Resnet
Stage Output Channels Kernel Size Resnet-101 Resnet-152
input 1280 - - -
stem 320 7 - -
@@ F.2 Resnet
classifier 2 1 - -
Table 5: ResNet-1D architectures used in ablation study.
We employed two variants of the ResNet-1D backbone in the ablation studies: ResNet-101 and ResNet-152. As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels. The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half. The difference lies in the number of bottleneck blocks per stage. ResNet1D-101 contains [3, 4, 23, 3] blocks across the four layers, while ResNet1D-152 is deeper with [3, 8, 36, 3] blocks. The final output is projected to the target classes through a $\times1$ convolutional layer.
Appendix G Data and Code Availability
The data and code are available at https://github.com/Jerry-alt19/BConformeR .

### v2

@@ 5 Conclusion and Limitation
There are three notable limitations to this study. First, all training and test interfaces are derived from in silico AlphaFold-Multimer models; although this choice enlarges the dataset, it may introduce misoriented chains or misplaced atoms. Second, antigen families and epitope types remain imbalanced, which could bias both training and evaluation. This is reflected in the significantly higher F1 scores observed on the 100-antigen test set compared with the 24-antigen blind test set, even though all records are unseen by BConformeR. Third, considering the computational cost, input sequences are limited to 1024 residues. Antigens longer than this threshold (e.g., 2nyy, 3ri5) cannot be directly processed by the model.
@@ Acknowledgments
Figure 6: Linear regions and discontinuous fragments. (A) (Left) The blue structure represents antigen 7c88, and the light green and light orange correspond to part of the antibody light chain and heavy chain, respectively; Linear regions are marked with red rectangles, and discontinuous fragments with green. (Right) Residues highlighted with numbers indicate epitopes. (B) A linear region of 7c88. Gap size of up to three non-contacting residues is chosen for a linear region on the basis of the structure of $\alpha$-helices.
@@ Appendix B Linear Regions and Discontinuous Fragments
Regions are defined as linear stretches of antigen sequence having at least three antibody-contacting residues (Fig. 6A). Gaps between contacting residues are allowed, and a gap size of up to three non-contacting residues is chosen on the basis of the structure of $\alpha$-helices, which enables residues on the same helical face to be included (Fig. 6B). Fragments, which do not contain enough antibody-contacting residues to qualify as a region (less than three), are considered discontinuous in our study (Fig. 6A). An algorithm classifying linear epitopes for an antigen sequence is shown below:
Algorithm 1 Linear Epitope Classification
1:label_tensorlabel\_tensor : sequence of residue labels
2:   ($1=$ contacting, $0=$ non-contacting, 100=-100= padding)
3:linear_flagslinear\_flags: boolean mask indicating linear epitope residues
4:Convert label_tensorlabel\_tensor to list labelslabels of length $L$
5:Initialize $linear\_flags\leftarrow[\text{False}]^{L}$, $visited\leftarrow[\text{False}]^{L}$
[…]
7:while $i<L$ do
8:  if $labels[i]\neq 1$ or $visited[i]=$ True then
9:   $i\leftarrow i+1$ $\triangleright$ skip non-epitope or already visited residue
10:   continue
11:  end if
12:  Initialize $segment\_indices\leftarrow[\,]$, $one\_count\leftarrow 0$, $zero\_count\leftarrow 0$
13:  $j\leftarrow i$
14:  while $j<L$ do
15:   if labels[j]=100labels[j]=-100 then
16:     break $\triangleright$ stop at padding
17:   else if $labels[j]=1$ then
18:     Append $j$ to segment_indicessegment\_indices
19:     $one\_count\leftarrow one\_count+1$, $zero\_count\leftarrow 0$
20:   else if $labels[j]=0$ then
21:     $zero\_count\leftarrow zero\_count+1$
22:     if zero_count>3zero\_count>3 then
23:      break $\triangleright$ gap too large
24:     end if
25:     Append $j$ to segment_indicessegment\_indices
26:   end if
27:   $j\leftarrow j+1$
28:  end while
29:  $ones\_in\_segment\leftarrow\{k\in segment\_indices\mid labels[k]=1\}$
30:  if $|ones\_in\_segment|\geq 3$ then
31:   for $k\in ones\_in\_segment$ do
32:     $linear\_flags[k]\leftarrow\text{True}$, $visited[k]\leftarrow\text{True}$
33:   end for
34:   $i\leftarrow segment\_indices[-1]+1$
35:  else
36:   $visited[i]\leftarrow\text{True}$, $i\leftarrow i+1$
37:  end if
38:end while
39:return linear_flagslinear\_flags
@@ C.1 Structure prediction
We initially obtained 1730 Ag-Ab FASTA files from AbEpiTope-1.0. However, due to the limited computational time and the large size of certain input sequences, we only processed 1300 complexes using the AlphaFold2 Multimer v3. Each of these 1300 complexes was modeled with 5 seeds, 6 models and 3 recycles, generating 30 structures per complex.
@@ C.2 Epitope statistics
The final dataset consists of 1180 Ag–Ab complexes, which contains a total of 1432 antigen chains and 2360 antibody chains. Epitopes were identified using a 4 Å distance cutoff between antigen and antibody residues. After classifying the epitopes into linear and discontinuous types, we found 2388 linear regions comprising 15,886 residues, and 2632 discontinuous fragments comprising 3574 residues.
@@ Appendix D Performance Metrics
The performance metrics used in this study include Precision, Recall, True Positive Rate (TPR), False Positive Rate (FPR), F1-score, Antigen-level intersection over union (AgIoU), Matthews correlation coefficient (MCC) and Pearson correlation coefficient (PCC), whose definitions are given as follows:
$\text{Precision}=\frac{TP}{TP+FP},$
@@
where $TP$, TNTN, FPFP and $FN$ denote the number of true positives, true negatives, false positives and false negatives, respectively; $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted values, while $\bar{y}$ and $\overline{\hat{y}}$ denote their means. Note that Recall and TPR are mathematically equivalent, but used in different contexts. The relationships between some of these metrics are given as:
$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}=\frac{2\cdot\text{AgIoU}}{1+\text{AgIoU}},$
@@ E.2 Baseline methods
Thresholds for all baselines were adjusted to reflect their best scoring behavior, ensuring a fair comparison with BConformeR in the blind test. We observed that SEPPA-3.0 performs better under a low decision threshold (around 0.1). For SEMA-1D 2.0, which defines epitope residues using an 8 Å distance cutoff, the predicted scores tend to be higher; thus, we adopted a threshold of 0.5 above its default of 0.351. Moreover, although the default threshold of BepiPred-3.0 is 0.1512, we found that a slightly higher value around 0.22 led to better balance between precision and recall on our test set.
@@ F.1 ViT1D
ViT1D-9 1536 1024 1 9 12 2.0
ViT1D-12 1536 1024 1 12 12 2.0
Table 4: ViT variants used in ablation study.
We conducted ablation studies on two variants of the ViT1D model differing in depth: ViT1D-9 and ViT1D-12. As shown in Table 4, both models share the same embedding dimension (1536), input sequence length (1024), patch size (1), number of attention heads (12), and MLP expansion ratio (2.0). The only difference is the number of transformer encoder layers.
F.2 ResNet1D
Stage Output Channels Kernel Size ResNet1D-101 ResNet1D-152
input 1280 - $\checkmark$ $\checkmark$
stem 320 7 $\checkmark$ $\checkmark$
@@ F.2 ResNet1D
classifier 2 1 $\checkmark$ $\checkmark$
Table 5: ResNet architectures used in ablation study.
We employed two variants of the ResNet1D backbone in the ablation studies: ResNet1D-101 and ResNet1D-152. As shown in Table 5, both models accept input feature maps with 1280 channels and begin with an initial convolutional layer producing 320 channels. The networks consist of four residual stages of bottlenecks with the expansion factor of 4, each stage starting by downsampling the channel dimension to half. The difference lies in the number of bottleneck blocks per stage. ResNet1D-101 contains [3, 4, 23, 3] blocks across the four layers, while ResNet1D-152 is deeper with [3, 8, 36, 3] blocks. The final output is projected to the target classes through a $\times1$ convolutional layer.